OpenStudy (anonymous):

Find the inverse f^-1(x) of the following function: f(x)=x^(2/3)+3x^(1/3)-2

OpenStudy (anonymous):

\[f(x)=x^{2/3}+3x ^{1/3}-2\]

OpenStudy (aum):

\[ \large f(x)=x^{2/3}+3x ^{1/3}-2 \\ \large f(x) = (x^{1/3})^2 + 3x^{1/3} + \frac 94 - \frac 94 - 2 \\ \large y = (x^{1/3} + \frac 32)^2 - \frac{17}{4} \\ \text{To find inverse, switch x and y and solve for y:} \\ \large x = (y^{1/3} + \frac 32)^2 - \frac{17}{4} \\ \large x + \frac{17}{4} = (y^{1/3} + \frac 32)^2 \\ \large \sqrt{x + \frac{17}{4} } = (y^{1/3} + \frac 32) \\ \large y^{1/3} = \left (\sqrt{x + \frac{17}{4} } - \frac 32 \right) \\ \large y = \left (\sqrt{x + \frac{17}{4} } - \frac 32 \right)^3 \\ \large f^{-1}(x) = \left (\sqrt{x + \frac{17}{4} } - \frac 32 \right)^3 \]

OpenStudy (aum):

or \[ \large f^{-1}(x) = \left (\frac{\sqrt{4x + 17}-3}{2} \right)^3 \]

OpenStudy (aum):

Inverse is valid for \(\large x \ge -\frac{17}{4}\)

OpenStudy (anonymous):

Thanks so much!